# Gatsby without the ‘E’: Creating Lipograms with LLMs

### v1

@@ 5 Lipogram Constraints vs Fidelity
To explore the malleability of written English, we employed our algorithm to create lipogram versions of Gatsby under constraints both stronger and weaker than forbidding the letter ‘E’. In particular, we generated lipogram translations for the first 200 paragraphs by excluding each single letter of the alphabet, plus selected subsets of vowels, consonants, and mixed letters to create even greater challenges. Representative paragraphs generated for each excluded letter of the alphabet appear in Appendix Table LABEL:table:avoided_characters.
@@
The results show a clear decay trend (clearer in Appendix A Figure 6): the cosine similarity decreases linearly with the exclusion fraction, indicating a steady loss of semantic fidelity even when removing up to 40% of the character space. Beyond this point similarity drops to noise levels, albeit greater than zero because our algorithm employs multi-selection. Excluding single vowels appears no more difficult than consonants of similar frequency, although excluding larger subsets of vowels becomes more crippling. Every English word must have a vowel: exclude all vowels and you only have numbers and punctuation left to express yourself with.
Sentences may remain semantically similar even as readability degrades. Figure 5 plots the corresponding increase in grammatical mistakes per paragraph in lipogram text as word selection gets progressively constrained. Grammatical fidelity largely remains intact up to the exclusion of ‘u’, at less than two mistakes per paragraph before rising sharply with greater restriction.
@@ 6 Conclusion
Lipogram translation demonstrates the power of modern language models, while highlighting the trade-offs between maintaining strict constraints and preserving text quality.
In exploring several LLMs we identified their respective strengths/weaknesses in handling constrained text generation. This study underscores the complexities of creating high-quality lipograms and provides a benchmark for future research in constrained text generation.
Left open for future work is the use of LLMs to produce coherent translations under even stronger palindromic constraints, where the text must read the same forward and backwards. Norvig (Norvig, 2016) used computer search techniques to produce a 21,012 word palindrome, but with no underlying story or meaning. Perhaps techniques akin to those developed in our work can be used produce interesting palindromic translations.

### v2

@@ 5 Lipogram Constraints vs Fidelity
To explore the malleability of written English, we employed our algorithm to create lipogram versions of Gatsby under constraints both stronger and weaker than forbidding the letter ‘e’. We generated lipogram translations for the first 200 paragraphs by excluding each single letter of the alphabet, plus selected subsets of vowels, consonants, and mixed letters to create even greater challenges.
@@
The results show a clear decay trend (clearer in Appendix A Figure 4): the cosine similarity decreases linearly with the exclusion fraction, indicating a steady loss of semantic fidelity even when removing up to 40% of the character space. Beyond this point similarity drops to noise levels. Excluding single vowels appears no more difficult than consonants of similar frequency, although excluding larger subsets of vowels becomes more crippling. Every English word must have a vowel: exclude all vowels and you only have numbers and punctuation left to express yourself with.
@@ 6 Conclusion
Lipogram translation demonstrates the power of modern language models, maintaining strict constraints while preserving text quality. In exploring several LLMs we identified their respective strengths/weaknesses in handling constrained text generation. This study underscores the complexities of creating high-quality lipograms and provides a benchmark for future research in constrained text generation.